OpenStudy (anonymous):

Whats Domain and Range for: h(x) = x^3 + 2x + 5?

OpenStudy (anonymous):

Is it just all real numbers for both? seems to easy...

OpenStudy (anonymous):

@tylerxy3 Is the equation currently in your calculator ?

OpenStudy (anonymous):

no?

OpenStudy (anonymous):

@tylerxy3 The domain is all real numbers correct !

OpenStudy (anonymous):

@tylerxy3 the range is infinite, because if you graph it you cannot find a min and max

OpenStudy (anonymous):

Do you just solve by graphing?

OpenStudy (anonymous):

so for range i put infinite?

OpenStudy (anonymous):

@tylerxy3 yes! You can even use https://www.desmos.com/calculator online

OpenStudy (anonymous):

So for f(x) = 1/x+3 I put Domain: x != 3 Range: all real numbers?

OpenStudy (anonymous):

or maybe that infinite to im not sure

OpenStudy (anonymous):

I also had g(x) = sqr(x+6) Domain:x>=-6 Range: all real numbers

OpenStudy (anonymous):

@tylerxy3 The domain is all real number its (-infinity symbol,infinity symbol)

OpenStudy (anonymous):

R = infinity symbl

OpenStudy (anonymous):

what about the other two?

OpenStudy (anonymous):

g(x) = sqr(x+6) ?

OpenStudy (anonymous):

fyi != means not equal thats my inner coder coming out sorry -_-

OpenStudy (anonymous):

squareroot of x+6

OpenStudy (anonymous):

Okay hang on doing it right now.. I'm also gunna see if I can give you a paper I have from my school that will help you. you won't ever have to graph just know the pattern

OpenStudy (anonymous):

oh sweet thank you

OpenStudy (anonymous):

a lot of this stuff is just harder online miss a lot of concepts

OpenStudy (anonymous):

Domain = [-6, infanti symbol)

OpenStudy (anonymous):

You use the bracket to show that its closed. and you use "(,)' so show it does on forever

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

to write range as infinite do u do range:[infinit symb]

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

okay did u see the f(x) = 1/x+3?

OpenStudy (anonymous):

I could be wrong but I think its

OpenStudy (anonymous):

(-infinity symbol 0] U [0,infinity symbol)

OpenStudy (anonymous):

(-infinity , symbol 0] U [0,infinity symbol) **

OpenStudy (anonymous):